# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 11.2.1 Fitted Q-Iteration
Fitted Q-Iteration (FQI, Ernst et al., 2005) is the simplest offline RL algorithm, predating the modern pessimism framework. FQI applies the standard Bellman backup iteratively using supervised regression on the fixed dataset, as described in Section 4. It does not include any explicit pessimism mechanism. When state-action coverage is poor, the max in the Bellman backup selects the highest Q-value among all actions at $s^{\prime}$, including actions never observed in the data. If the function approximator generalizes poorly at these unseen actions, targets become noisy and biased upward, causing the overestimation cascade. FQI works well when coverage is good but degrades as coverage gaps grow.196196196Munos and Szepesvári (2008) prove finite-time error bounds for FQI under approximate Bellman completeness and all-policy concentrability. Both assumptions are strong, and violation of either leads to divergence in practice.
@@ 11.2.4 Batch-Constrained Q-Learning
11.3 Simulation: Offline RL for Dynamic Pricing
The simulation study evaluates the four offline RL algorithms on a perishable inventory pricing problem with demand regime switching. A retailer with $I_{\max}=30$ units of perishable inventory must set prices over $H=20$ periods. The state $(i,d,t)$ consists of current inventory $i\in\{0,\ldots,30\}$, demand regime $d\in\{1,2,3,4\}$, and time remaining $t\in\{1,\ldots,20\}$. The action is a price $p\in\{1,\ldots,10\}$. Demand follows $Q\sim\text{Poisson}(\lambda_{0}[d]\cdot e^{-0.15p})$ with base rates $\lambda_{0}=(1.5,3.0,5.0,8.0)$, and the reward is $r=p\cdot\min(Q,i)$. Demand regimes follow a 4-state Markov chain with diagonal persistence 0.6. Unsold inventory at the terminal period incurs a spoilage cost of \$2.00 per unit, making clearance pricing valuable near the deadline.197197197The spoilage penalty creates distributional shift. The optimal policy adapts prices to inventory level and time remaining, using lower prices near the deadline when inventory is high. The behavioral policy ignores these state variables and prices at the maximum, so the Q-function at state-adapted pricing actions is extrapolation from sparse data. The \$2.00 penalty is a deliberate design choice: under harsher penalties (e.g., \$10 per unit), all methods collapse to 48–53% of optimal regardless of algorithmic sophistication, confirming that no offline correction can overcome severe distributional shift when the penalty regime amplifies consequences of the behavioral policy’s suboptimality. The behavioral policy represents a conservative pricing team that always sets the maximum price ($p=10$) regardless of demand regime, inventory, or time remaining, with probability 0.85 and randomizes uniformly over all prices with probability 0.15.198198198With 500 episodes (10,000 transitions) on a state-action space of 24,800 pairs, the 85% concentration at price 10 ensures that the behavioral state-action occupancy diverges significantly from the optimal policy’s occupancy, while the 15% uniform component provides sparse off-policy coverage. All episodes start at full inventory ($i=30$). All offline methods train on 500 episodes and are evaluated over 1,000 episodes against the DP optimal policy computed by backward induction.199199199FQI uses the standard Bellman backup with $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$, deliberately without a target network, to isolate the overestimation cascade as a pedagogical baseline. Adding target networks to FQI mitigates but does not eliminate extrapolation error. CQL and IQL include target networks following their original implementations (Kumar et al., 2020; Kostrikov et al., 2022); for CQL, target networks proved essential, as the conservative penalty amplifies bootstrap instability without them. Results are averaged over 20 independent seeds.
@@ 11.3 Simulation: Offline RL for Dynamic Pricing
FQI achieves 81.2% of the DP optimal, substantially below the behavioral cloning baseline of 88.0% (Table 20). Without any mechanism to control extrapolation error, the $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ operator in FQI’s Bellman backup selects overestimated Q-values at out-of-distribution actions, and these overestimates compound across 200 iterations of fitted Q-iteration. CQL and IQL both exceed the behavioral baseline, achieving 91.9% and 92.0% respectively.200200200CQL uses $\alpha=0.1$, the result of a search over $\alpha\in\{5.0,2.0,0.5,0.1\}$. Larger values push Q-values down too aggressively, collapsing the learned policy to the behavioral action at most states; the right $\alpha$ is problem-specific and can vary by orders of magnitude. CQL’s conservative penalty suppresses Q-values at actions not well-represented in the data while preserving relative ordering among data-supported actions, allowing the policy to discover lower prices that clear inventory near the deadline. IQL achieves the same improvement through a different mechanism: by replacing $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ with the expectile-regressed value function $V(s^{\prime})$ as the Bellman continuation, IQL avoids querying Q at unseen actions during training while still extracting a policy that improves on the behavioral at states where the 15% noise component revealed better pricing actions. BCQ matches the behavioral at 88.0%; its action constraint restricts the policy to prices near 10, preventing both overestimation and improvement.201201201When the behavioral policy concentrates 85% probability at a single action, BCQ’s threshold constraint permits only that action at most states, effectively reducing BCQ to behavioral cloning regardless of the learned Q-values.
@@
Figure 20 varies the behavioral noise parameter $\epsilon_{b}\in\{0.05,0.3,0.9\}$. CQL and IQL remain above the BC baseline across all coverage levels, confirming that both pessimism mechanisms provide robustness to the data distribution. FQI exhibits non-monotone behavior: it peaks at moderate coverage ($\epsilon_{b}=0.3$) where partial exploration controls the overestimation cascade, but collapses at both extremes.202202202At high $\epsilon_{b}$, the near-uniform behavioral policy provides Q-function targets across all actions, giving the unconstrained $\max_{a^{\prime}}$ operator more opportunities to select overestimated values rather than fewer. BCQ collapses at $\epsilon_{b}=0.9$ because a nearly uniform behavioral policy renders the action constraint vacuous, reducing BCQ to unconstrained FQI. These coverage patterns connect directly to the concentrability framework (Definition D2): as $\epsilon_{b}$ decreases, the concentrability coefficient $C^{*}$ grows because the optimal policy’s state-action occupancy diverges from the behavioral policy’s, and the $1/\sqrt{N_{h}(s_{h},a_{h})}$ terms in (98) become large at states the optimal policy visits. CQL and IQL remain robust because their conservative adjustments scale with data sparsity, while FQI lacks this adaptive correction.
@@ 11.5 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).203203203Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy204204204An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (104) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 11.6 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)205205205Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$,206206206After fine-tuning concludes, the resulting LLM is deployed with frozen weights. Each user interaction is a forward pass in the execution phase (Section 2); the model does not update its parameters from conversations. Periodic retraining on new preference data constitutes a separate training phase. with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@ 11.7 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.208208208The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).209209209The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
@@ 11.8 Simulation Study: Preference Learning in Job Search
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.210210210The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (104) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.211211211DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 108) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
@@
Figure 23 reports the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.212212212DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.213213213DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
@@
Table 21 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.214214214DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.215215215An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
@@
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that economists typically have access to.216216216This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.

### 23 Mar 2026 v4

@@ 11.2.1 Fitted Q-Iteration
Fitted Q-Iteration (FQI, Ernst et al., 2005) is the simplest offline RL algorithm, predating the modern pessimism framework. FQI applies the standard Bellman backup iteratively using supervised regression on the fixed dataset, as described in Section 4. It does not include any explicit pessimism mechanism. When state-action coverage is poor, the max in the Bellman backup selects the highest Q-value among all actions at $s^{\prime}$, including actions never observed in the data. If the function approximator generalizes poorly at these unseen actions, targets become noisy and biased upward, causing the overestimation cascade. FQI works well when coverage is good but degrades as coverage gaps grow.206206206Munos and Szepesvári (2008) prove finite-time error bounds for FQI under approximate Bellman completeness and all-policy concentrability. Both assumptions are strong, and violation of either leads to divergence in practice.
@@ 11.2.4 Batch-Constrained Q-Learning
11.3 Simulation Study: Offline RL for Dynamic Pricing
The simulation study evaluates the four offline RL algorithms on a perishable inventory pricing problem with demand regime switching. A retailer with $I_{\max}=30$ units of perishable inventory must set prices over $H=20$ periods. The state $(i,d,t)$ consists of current inventory $i\in\{0,\ldots,30\}$, demand regime $d\in\{1,2,3,4\}$, and time remaining $t\in\{1,\ldots,20\}$. The action is a price $p\in\{1,\ldots,10\}$. Demand follows $Q\sim\text{Poisson}(\lambda_{0}[d]\cdot e^{-0.15p})$ with base rates $\lambda_{0}=(1.5,3.0,5.0,8.0)$, and the reward is $r=p\cdot\min(Q,i)$. Demand regimes follow a 4-state Markov chain with diagonal persistence 0.6. Unsold inventory at the terminal period incurs a spoilage cost of \$2.00 per unit, making clearance pricing valuable near the deadline.207207207The spoilage penalty creates distributional shift. The optimal policy adapts prices to inventory level and time remaining, using lower prices near the deadline when inventory is high. The behavioral policy ignores these state variables and prices at the maximum, so the Q-function at state-adapted pricing actions is extrapolation from sparse data. The \$2.00 penalty is a deliberate design choice: under harsher penalties (e.g., \$10 per unit), all methods collapse to 48–53% of optimal regardless of algorithmic sophistication, confirming that no offline correction can overcome severe distributional shift when the penalty regime amplifies consequences of the behavioral policy’s suboptimality. The behavioral policy represents a conservative pricing team that always sets the maximum price ($p=10$) regardless of demand regime, inventory, or time remaining, with probability 0.85 and randomizes uniformly over all prices with probability 0.15.208208208With 500 episodes (10,000 transitions) on a state-action space of 24,800 pairs, the 85% concentration at price 10 ensures that the behavioral state-action occupancy diverges significantly from the optimal policy’s occupancy, while the 15% uniform component provides sparse off-policy coverage. All episodes start at full inventory ($i=30$). All offline methods train on 500 episodes and are evaluated over 1,000 episodes against the DP optimal policy computed by backward induction.209209209FQI uses the standard Bellman backup with $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$, deliberately without a target network, to isolate the overestimation cascade as a pedagogical baseline. Adding target networks to FQI mitigates but does not eliminate extrapolation error. CQL and IQL include target networks following their original implementations (Kumar et al., 2020; Kostrikov et al., 2022); for CQL, target networks proved essential, as the conservative penalty amplifies bootstrap instability without them. Results are averaged over 20 independent seeds.
@@ 11.3 Simulation Study: Offline RL for Dynamic Pricing
FQI achieves 81.2% of the DP optimal, substantially below the behavioral cloning baseline of 88.0% (Table 20). Without any mechanism to control extrapolation error, the $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ operator in FQI’s Bellman backup selects overestimated Q-values at out-of-distribution actions, and these overestimates compound across 200 iterations of fitted Q-iteration. CQL and IQL both exceed the behavioral baseline, achieving 91.9% and 92.0% respectively.210210210CQL uses $\alpha=0.1$, the result of a search over $\alpha\in\{5.0,2.0,0.5,0.1\}$. Larger values push Q-values down too aggressively, collapsing the learned policy to the behavioral action at most states; the right $\alpha$ is problem-specific and can vary by orders of magnitude. CQL’s conservative penalty suppresses Q-values at actions not well-represented in the data while preserving relative ordering among data-supported actions, allowing the policy to discover lower prices that clear inventory near the deadline. IQL achieves the same improvement through a different mechanism: by replacing $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ with the expectile-regressed value function $V(s^{\prime})$ as the Bellman continuation, IQL avoids querying Q at unseen actions during training while still extracting a policy that improves on the behavioral at states where the 15% noise component revealed better pricing actions. BCQ matches the behavioral at 88.0%; its action constraint restricts the policy to prices near 10, preventing both overestimation and improvement.211211211When the behavioral policy concentrates 85% probability at a single action, BCQ’s threshold constraint permits only that action at most states, effectively reducing BCQ to behavioral cloning regardless of the learned Q-values.
@@
Figure 21 varies the behavioral noise parameter $\epsilon_{b}\in\{0.05,0.3,0.9\}$. CQL and IQL remain above the BC baseline across all coverage levels, confirming that both pessimism mechanisms provide robustness to the data distribution. FQI exhibits non-monotone behavior: it peaks at moderate coverage ($\epsilon_{b}=0.3$) where partial exploration controls the overestimation cascade, but collapses at both extremes.212212212At high $\epsilon_{b}$, the near-uniform behavioral policy provides Q-function targets across all actions, giving the unconstrained $\max_{a^{\prime}}$ operator more opportunities to select overestimated values rather than fewer. BCQ collapses at $\epsilon_{b}=0.9$ because a nearly uniform behavioral policy renders the action constraint vacuous, reducing BCQ to unconstrained FQI. These coverage patterns connect directly to the concentrability framework (Definition D2): as $\epsilon_{b}$ decreases, the concentrability coefficient $C^{*}$ grows because the optimal policy’s state-action occupancy diverges from the behavioral policy’s, and the $1/\sqrt{N_{h}(s_{h},a_{h})}$ terms in (103) become large at states the optimal policy visits. CQL and IQL remain robust because their conservative adjustments scale with data sparsity, while FQI lacks this adaptive correction.
@@ 11.5 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).213213213Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy214214214An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (109) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 11.6 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)215215215Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$,216216216After fine-tuning concludes, the resulting LLM is deployed with frozen weights. Each user interaction is a forward pass in the execution phase (Section 2); the model does not update its parameters from conversations. Periodic retraining on new preference data constitutes a separate training phase. with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@ 11.7 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.218218218The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).219219219The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
@@ 11.8 Simulation Study: Preference Learning in Job Search
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.220220220The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (109) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.221221221DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 113) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
@@
Figure 24 reports the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.222222222DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.223223223DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
@@
Table 21 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.224224224DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.225225225An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
@@
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that structural modelers typically have access to.226226226This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.